# Sublinear bilipschitz equivalence and the quasiisometric classification of solvable Lie groups

### v2

@@ 2.C. SBE Preserve Product Structure
First we show there is $d_{0}(x)$ such that for $d>d_{0}(x)$ there are no $d$-separated $i$-horizontal pairs in $B(x,99d)$ that are $j$-semi-compressed. Assume towards contradiction there were such pairs for all $d>0$. They give rise to a sequence of radii $d_{k}\rightarrow\infty$ and points $x^{k},z^{k}\in B(x,99d_{k})$ that are $i$-horizontal, $d(x^{k},z^{k})=d_{k}$, with $(x^{k},z^{k})$ being $j$-semi-compressed. Consider the asymptotic cone $X_{\omega}:=\operatorname{Cone}_{\omega}(X,x,d_{k})$, and the points $x_{\omega}:=(x^{k})_{k\in\mathbf{N}},z_{\omega}:=(z^{k})_{k\in\mathbf{N}}\in X_{\omega}$. These points agree on the $(X_{\omega})_{m}$-coordinate for all $m\neq i$. The hypothesis on $X$ and $Y$ assures that the cone map $F=\operatorname{Cone}(\phi)$ respects the product structure of $X_{\omega}$ and $Y_{\omega}$, therefore if $m\neq i$ we have
(2.1) $d_{(Y_{\omega})_{m}}\Big{(}\pi_{(Y_{\omega})_{m}}\big{(}F(x_{\omega})\big{)},\pi_{(Y_{\omega})_{m}}\big{(}F(z_{\omega})\big{)}\Big{)}=0$
and so
(2.2) $d_{(Y_{\omega})}\big{(}F(x_{\omega}),F(z_{\omega})\big{)}=d_{(Y_{\omega})_{i}}\Big{(}\pi_{(Y_{\omega})_{i}}\big{(}F(x_{\omega})\big{)},\pi_{(Y_{\omega})_{i}}\big{(}F(z_{\omega})\big{)}\Big{)}.$
Finally, for every $m\in\{1,\dots,n\}$ it holds that $\pi_{(Y_{\omega})_{m}}\big{(}F(x_{\omega})\big{)}=\lim_{\omega}\pi_{Y_{m}}\big{(}\phi(x^{k})\big{)}$ and similarly for $z_{\omega}$, yielding
$\lim_{\omega}\frac{1}{d_{k}}d_{Y_{m}}\Big{(}\pi_{Y_{m}}\big{(}\phi(x^{k})\big{)},\pi_{Y_{m}}\big{(}\phi(z^{k})\big{)}\Big{)}=d_{(Y_{\omega})_{m}}\Big{(}\pi_{(Y_{\omega})_{m}}\big{(}F(x_{\omega})\big{)},\pi_{(Y_{\omega})_{m}}\big{(}F(z_{\omega})\big{)}\Big{)}.$
Plugging this formula into Equations 2.1 and 2.2 above, we get:
@@
    For $j\neq i$ $\lim_{\omega}\frac{1}{d_{k}}d_{Y_{j}}\Big{(}\pi_{Y_{j}}\big{(}\phi(x^{k})\big{)},\pi_{Y_{j}}\big{(}\phi(z^{k})\big{)}\Big{)}=0$.
@@
    For $j=i$, the fact that $F$ is $L$-bilipschitz and preserves the product structure, together with $d_{X_{\omega}}(x_{\omega},z_{\omega})=1$, implies
    $\lim_{\omega}\frac{1}{d_{k}}d_{Y_{i}}\Big{(}\pi_{Y_{i}}\big{(}\phi(x^{k})\big{)},\pi_{Y_{i}}\big{(}\phi(z^{k})\big{)}\Big{)}\in[L^{-1},L].$
The assumption that $0<\epsilon<L_{1}^{-1}<L^{-1}$ contradicts the assumption that $(x^{k},z^{k})$ are $j$-semi-compressed.
Next, we show that there is $d_{0}(x)$ such that for every $d>0$, every two pairs of $d$-separated $i$-horizontal points inside $B(x,99d)$ are $j$-compatible. This is done exactly as above. If there is no such $d_{0}(x)$, we obtain sequences of radii $d_{k}\to\infty$ and points $x^{1,k},z^{1,k},x^{2,k},z^{2,k}\in B(x,99d_{k})$ such that for each $k\in\mathbf{N}$ and $l\in\{1,2\}$, the pair $(x^{l,k},z^{l,k})$ is $i$-horizontal with $d_{X}(x^{l,k},z^{l,k})\in[d_{k},99d_{k})$ and $(x^{1,k},z^{1,k})$ is $j$-compressed while $(x^{2,k},z^{2,k})$ is $j$-uncompressed. We consider the cones $X_{\omega}:=\operatorname{Cone}(X,x,d_{k}),Y_{\omega}:=\operatorname{Cone}\big{(}Y,\phi(x),d_{k}\big{)}$ and the cone map $F:=\operatorname{Cone}(\phi)$. For $l\in\{1,2\}$, the points $x_{\omega}^{l}=(x^{l,k})_{k\in\mathbf{N}},z_{\omega}^{l}=(z^{l,k})_{k\in\mathbf{N}}\in X_{\omega}$ both lie within the ball of radius 99 around the base point of the cone, and agree on the $m\neq i$ coordinates. Since $F$ respects products there are $L$-bilipschitz homeomorphisms $F_{m}:(X_{\omega})_{m}\rightarrow(Y_{\omega})_{m}$ for which $F_{i}\circ\pi_{(X_{\omega})_{m}}=\pi_{(Y_{\omega})_{m}}\circ F$, hence for $m\neq i$ we must have for $l\in\{1,2\}$
$d_{(Y_{\omega})_{m}}\Big{(}\pi_{(Y_{\omega})_{m}}\big{(}F(x_{\omega}^{l})\big{)},\pi_{(Y_{\omega})_{m}}\big{(}F(z_{\omega}^{l})\big{)}\Big{)}=0$
Since this does not depend on $l$, we see as in step 1 that for $j\neq i$ for all large enough $k$, both $(x^{1,k},z^{1,k})$ and $(x^{2,k},z^{2,k})$ are $j$-compressed and in particular compatible. A similar argument shows that these pairs are both $j$-uncompressed for $j=i$. We conclude that the pairs are $j$-compatible for all $j$.
Step 22
We now show that we can choose $d_{0}$ to be a sublinear function in $O(u)$. We start by taking, for each $x\in X$, the infimal $\tilde{d}_{0}(x)$ such that for all $d\geq\tilde{d}_{0}(x)$, all $i$-horizontal $d$-separated pairs in $B(x,99d)$ are $j$-compatible. Next define the radial function $d_{0}(|x|):=\sup_{y:|y|=|x|}\tilde{d}_{0}(y)$.
A-priori $d_{0}:\mathbf{R}_{\geq 0}\rightarrow\mathbf{R}_{\geq 0}\cup\{\infty\}$, and we need to show it is in fact real valued. Assume towards contradiction that $d_{0}$ is not bounded along some sphere $S$ around $x_{0}$, and get points $x^{k}\in S$ and real numbers $r_{k}\rightarrow\infty$ such that the balls $B(x^{k},10r_{k})$ contain two $i$-horizontal $r_{k}$-separated pairs $(x^{1,k},z^{1,k}),(x^{2,k},z^{2,k})$ that are $j$-incompatible. We consider the cone $X_{\omega}:=(X,x_{0},r_{k})$, based at $x_{\omega}=(x_{0})$. Since $x^{k}\in S$ we have $d(x^{k},x_{0})=R$ for some constant $R$, so $(x^{l,k}),(z^{l,k})\in X_{\omega}$. Moreover, the assumption that $(x^{1,k},z^{1,k})$ are $r_{k}$-separated and lie in $B(x^{k},99r_{k})$ means that $d_{\omega}(x_{\omega}^{l},z_{\omega}^{l})\in[1,198]$. Again, for every $k$ $x^{l,k}$ and $z^{l,k}$ agree on the $m\neq i$ coordinates, hence so do $x^{l}_{\omega}$ and $z^{l}_{\omega}$. As in previous steps, we get a contradiction to the assumption that $(x^{1,k},z^{1,k}),(x^{2,k},z^{2,k})$ are $j$-incompatible.
The function $d_{0}$ is therefore a radial real valued function. We show it is $O(u)$. Assume towards contradiction that there is a sequence $x^{k}\rightarrow\infty$ so that $\lim_{\omega}\frac{u(x^{k})}{d_{0}(x^{k})}=0$. By the construction of $d_{0}$ we may take $x^{k}$ such that each ball $B\big{(}x^{k},99d_{0}(x^{k})\big{)}$ contains two $(d_{0}(x^{k})-1)$-separated $i$-horizontal pairs that are $j$-incompatible. The assumption $\lim_{\omega}\frac{u(x^{k})}{d_{0}(x^{k})}=0$ implies that the cone with moving base points $X_{\omega}:=\operatorname{Cone}_{\omega}\big{(}X,x^{k},d_{0}(x^{k})\big{)}$ is $u$-admissible. By Lemma 2.7 the cone map $F=\operatorname{Cone}(\phi)$ is $L$-bilipschitz, and so respects products. One gets a contradiction as in step $2$.
Step 33
We have established the following situation: there is $d_{0}\in O(u)$ such that for every $x\in X$ and every $d\geq d_{0}(x)$, the property of whether a $d$-separated $i$-horizontal pair inside $B=B(x,99d)$ is $j$-uncompressed or $j$-compressed depends only on $x$, $i$, $j$ and $d$, and not on the specific pair (in particular such a pair cannot be $j$-semi-compressed). We call it the $j$-type of $B$. We assume, as we may, that $d_{0}$ is subadditive. We can right away conclude that there are no $d_{0}$-separated $i$-horizontal $j$-semi-compressed pairs at all in $X$: such a pair $(x,x^{\prime})$ must admit $d(x,x^{\prime})\geq d_{0}(x)$ and therefore for $d=d(x,x^{\prime})\geq d_{0}(x)$, $x,x^{\prime}$ are $d$-separated and $x^{\prime}\in B(x,99d)$. Step $1$ shows that $(x,x^{\prime})$ is not $j$-semi-compressed.
Our next goal is to lose any restriction on the radius of the balls $B(x,99d)$. Fix $x\in X$. Our first observation is that the $j$-type of $d_{0}$-separated $i$-horizontal pairs which involve $x$ depend only on $x$: they are the type of the ball $B:=B\big{(}x,99d_{0}(x)\big{)}$. Indeed if $(x,x^{\prime})$ is an $i$-horizontal pair, one can consider the geodesic $\gamma:=[x,x^{\prime}]$. If $x^{\prime}\in B$ the claim is immediate. In particular the pair $\Big{(}x,\gamma\big{(}98d_{0}(x)\big{)}\Big{)}$ has the $j$-type of $B$. Therefore for $R_{1}:=98d_{0}(x)$, the ball $B_{1}:=B(x,99\cdot R_{1})$ has the $j$-type of $B$. We can continue to enlarge the radius of the balls until we find one which contains $x^{\prime}$, so indeed $(x,x^{\prime})$ has the $j$-type of $B$, which is what we were set out to prove.
We push the above argument a bit further in order to show that any two $d_{0}$-separated $i$-horizontal pairs are $j$-compatible, regardless of which leaf of $X_{i}$ they are in. Let $(x,x^{\prime})$ and $(z,z^{\prime})$ be two $d_{0}$-separated $i$-horizontal pairs. Assume first that $x,x^{\prime},z,z^{\prime}\in X$ are pairwise $i$-horizontal, so they all lie in the same leaf. Since each $X_{i}$ can be assumed to have infinite diameter, we can find a point $p$ in the same $X_{i}$ leaf which is far enough so that $(p,w)$ is $d_{0}$-separated for all $w\in\{x,x^{\prime},z,z^{\prime}\}$. It is clear that the $j$-type of all points must be equal to that of $B\big{(}p,99d_{0}(p)\big{)}$.
This allows us to define the $j$-type of each $X_{i}$-leaf, defined by a choice of $\hat{x}\in\hat{X}:=\prod_{m\neq i}X_{m}$. To finish the proof we need to show that the $(i,j)$-type of $\hat{x}\in\hat{X}$ depends only on $(i,j)$ and not on the choice of $\hat{x}$. Indeed, fix $\hat{x},\hat{z}\in\prod_{m\neq i}X_{m}$, and $\gamma:\mathbf{R}_{\geq 0}\rightarrow X_{i}$ an infinite ray in $X_{i}$ (we assume $X_{i}$ is of infinite diameter and geodesic and so by an Arzela-Ascoli argument there exists an infinite geodesic ray).
Consider the points $x^{k}=\big{(}\gamma(k),\hat{x}\big{)},z^{k}=\big{(}\gamma(k),\hat{z}\big{)}\in X$. For all large enough $k$, $(x^{0},x^{k})$ and $(z^{0},z^{k})$ are $d_{0}$-separated $i$-horizontal. For all large enough $k,m\in\mathbb{N}$, all pairs $(x^{0},x^{k}),(x^{0},x^{m})$ are $j$-compatible and all $(z^{0},z^{k}),(z^{0},z^{m})$ are $j$-compatible. Consider the cone $X_{\omega}:=(X,x^{0},k)$. Let $x_{\omega}^{0}$ be the base point of $X_{\omega}$, $x_{\omega}:=(x^{k})_{k},z_{\omega}:=(z^{k})_{k}$. Clearly $x_{\omega}=z_{\omega}$ and $d(x_{\omega}^{0},x_{\omega})=1$. We conclude that for all large enough $k,m\in\mathbb{N}$, it holds that $(x^{0},x^{k})$ and $(z^{0},z^{m})$ are $j$-compatible, so $\hat{x},\hat{z}$ have the same $(i,j)$-type. The lemma stands proven. ∎
Proof of Proposition 2.11.
@@
Together with Lemma 2.17, the assumption on the product decomposition of the relevant cone maps implies that for every $i$ there is a unique $j$ for which the projection on $Y_{j}$ is uncompressed for all $i$-horizontal pairs. We may reindex and assume $Y_{i}$ is that factor. To ease the notation, we fix from now $i=1$. The proof is identical for all other $i$.
@@
$L_{1}^{-1}d(z,z^{\prime})\leqslant d\Big{(}\pi_{Y_{1}}\big{(}\phi(z,\hat{x})\big{)},\pi_{Y_{1}}\big{(}\phi(z^{\prime},\hat{x})\big{)}\Big{)}\leqslant L_{1}d(z,z^{\prime})$

### v3

@@ 2.C. Proof of Theorem C
In the notations of the statement, the hypothesis $\lim_{\omega}\frac{1}{\sigma_{k}}d\big(\pi_{i}(z_{k}),\pi_{i}(z_{k}^{\prime})\big)=0$ implies $\pi_{\omega_{i}}(z_{\omega})=\pi_{\omega_{i}}(z_{\omega}^{\prime})$, where $\pi_{\omega_{i}}$ is the projection to the $i$-th factor in the cone. Therefore the assumption that $\phi_{\omega}$ preserves the product structure gives $\pi_{\omega_{i}}\big(\phi_{\omega}(z_{\omega})\big)=\pi_{\omega_{i}}\big(\phi_{\omega}(z_{\omega}^{\prime})\big)$. Item (1) follows. Under the assumption of (2), we thus have
$d_{Y\omega}\big(\phi_{\omega}(z_{\omega}),\phi_{\omega}(z_{\omega}^{\prime})\big)=d_{Y_{\omega_{i}}}\Big(\pi_{\omega_{i}}\big(\phi_{\omega}(z_{\omega})\big),\pi_{\omega_{i}}\big(\phi_{\omega}(z_{\omega}^{\prime})\big)\Big).$
Since $X_{\omega}$ is $u$-admissible, $\phi_{\omega}$ is $L$-bilipschitz by Lemma 2.8, hence
(2.1) $\frac{1}{L}d_{\omega}(z_{\omega},z_{\omega}^{\prime})\leq d_{Y_{\omega_{1}}}\Big(\pi_{i}\big(\phi(z_{\omega})\big),\pi_{i}\big(\phi(z_{\omega}^{\prime})\big)\Big)\leq Ld_{\omega}(z_{\omega},z_{\omega}^{\prime}).$
By the definition of the distance in the cone, we have
$\lim_{\omega}\frac{d_{Y_{i}}\Big(\pi_{i}\big(\phi(z_{k})\big),\pi_{i}\big(\phi(z_{k}^{\prime})\big)\Big)}{d_{X}(z_{k},z_{k}^{\prime})}$ $=\frac{\lim_{\omega}\frac{1}{\sigma_{k}}d_{Y_{i}}\Big(\pi_{i}\big(\phi(z_{k})\big),\pi_{i}\big(\phi(z_{k}^{\prime})\big)\Big)}{\lim_{\omega}\frac{1}{\sigma_{k}}d_{X}(z_{k},z_{k}^{\prime})}$
$=\frac{d_{Y_{\omega_{1}}}\Big(\pi_{i}\big(\phi(z_{\omega})\big),\pi_{i}\big(\phi(z_{\omega}^{\prime})\big)\Big)}{d_{\omega}(z_{\omega},z_{\omega}^{\prime})}\in\left[\frac{1}{L},L\right],$
which proves the second item. Assume further that the pairs $(z_{k},z_{k}^{\prime})$ defining $z_{\omega},z_{\omega}^{\prime}\in X_{\omega}$ are $i$-horizontal $\omega$-almost surely. In particular, the assumption of (2) holds (and automatically the hypothesis of (1) for $j\neq i$). Since $L_{1}>L$, this immediately implies that $\omega$-almost surely the pairs $(z_{k},z_{k}^{\prime})$ are $i$-uncompressed. Fix $j\neq i$. By definition of the cone, $z_{\omega},z_{\omega}^{\prime}\in X_{\omega}$ implies $\lim_{\omega}\frac{d_{X}(z_{k},z_{k}^{\prime})}{\sigma_{k}}=C$ for some $C<\infty$, so there are $\epsilon_{k}\geq 0$ such that $C\sigma_{k}-\epsilon_{k}\leq d_{X}(z_{k},z_{k}^{\prime})$ and $\lim_{\omega}\frac{\epsilon_{k}}{\sigma_{k}}=0$. Therefore $\omega$-almost surely
$\sigma_{k}\frac{C}{2}<\sigma_{k}(C-\frac{\epsilon_{k}}{\sigma_{k}})=\sigma_{k}C-\epsilon_{k}\leq d_{X}(z_{k},z_{k}^{\prime}).$
We then have that $\omega$-almost surely
$\frac{d_{Y_{j}}\Big(\pi_{j}\big(\phi(z_{k})\big),\pi_{j}\big(\phi(z_{k}^{\prime})\big)\Big)}{d_{X}(z_{k},z_{k}^{\prime})}\leq\frac{d_{Y_{j}}\Big(\pi_{j}\big(\phi(z_{k})\big),\pi_{j}\big(\phi(z_{k}^{\prime})\big)\Big)}{\sigma_{k}\frac{C}{2}}.$
The ultralimit of the right-hand side is 0 by the conclusion of the first item of the lemma. Therefore for every $\epsilon$, $\omega$-almost surely the left hand side is smaller than $\epsilon$. We conclude that $\omega$-almost surely $(z_{k},z_{k}^{\prime})$ is $j$-compressed for all $j\neq i$. ∎
The following lemma is the core of the proof of Theorem C.
Lemma 2.16 ($O(u)$ version of Lemma $2.7$ in [KKL98]).
There exists a subadditive nondecreasing sublinear function $d_{0}\in O(u)$ such that for every fixed $i,j\in\{1,\dots,n\}$, either all $d_{0}$-separated $i$-horizontal pairs are $j$-compressed or all such pairs are $j$-uncompressed.
Remark 2.17.
Parts of the proof below (specifically in step 2) could have been made slightly simpler under the assumption that the $X_{i}$ and $Y_{j}$ are uniformly discrete. While natural in our context, we do not impose this assumption here.
Proof.
[…]
@@
First we show there is $d_{0}(x)$ such that for $d\geq d_{0}(x)$ there are no $d$-separated $i$-horizontal pairs in $B(x,99d)$ that are $j$-semi-compressed. Assume towards contradiction there were such pairs for all $d>0$. These pairs give rise to a sequence of radii $d_{k}\rightarrow\infty$ and points $x_{k},z_{k}\in B(x,99d_{k})$ that are $i$-horizontal, $d(x_{k},z_{k})=d_{k}$, with $(x_{k},z_{k})$ being $j$-semi-compressed. Consider the asymptotic cone $X_{\omega}:=\operatorname{Cone}_{\omega}(X,x,d_{k})$. It is $u$-admissible since the base point $x$ is fixed. Moreover, $x_{k},z_{k}\in B(x,99d_{k})$ so we have $x_{\omega}:=[(x_{k})_{k}],z_{\omega}:=[(z_{k})_{k}]\in X_{\omega}$. Since the pairs are $i$-horizontal, Lemma 2.15 yields a contradiction: $\omega$-almost surely, $(x_{k},z_{k})$ is $l$-compressed for $l\neq i$ and $i$-uncompressed. We conclude that there is $d_{0}(x)$ such that $d>d_{0}(x)$ implies that inside $B(x,99d)$ there are no $i$-horizontal $d$-separated $j$-semi-compressed pairs.
With a similar argument, we now show that there is $d_{0}(x)$ such that for every $d\geq d_{0}(x)$, every two pairs of $d$-separated $i$-horizontal points inside $B(x,99d)$ are $j$-compatible. Assume towards contradiction there is no such $d_{0}(x)$. One obtains a sequence of radii $d_{k}\to\infty$ and sequences $x_{1,k},z_{1,k},x_{2,k},z_{2,k}\in B(x,99d_{k})$ such that for all $k$ and $\alpha\in\{1,2\}$, the pair $(x_{\alpha,k},z_{\alpha,k})$ is $i$-horizontal with $d_{X}(x_{\alpha,k},z_{\alpha,k})\in[d_{k},99d_{k})$ and $(x_{1,k},z_{1,k})$ is $j$-compressed while $(x_{2,k},z_{2,k})$ is $j$-uncompressed. We consider the $u$-admissible cone $X_{\omega}:=\operatorname{Cone}_{\omega}(X,x,d_{k})$. Since $d_{X}(x_{\alpha,k},z_{\alpha,k})\in[d_{k},99d_{k})$, we see that $x_{\omega,\alpha}:=[(x_{\alpha,k})_{k}],z_{\omega,\alpha}:=[(z_{\alpha,k})_{k}]\in X_{\omega}$. Since all pairs are $i$-horizontal, we conclude from Lemma 2.15 that if $l\neq i$, $\omega$-almost surely the pair $(x_{2,k},z_{2,k})$ is $l$-compressed. The assumption on $(x_{2,k},z_{2,k})$ forces $j=i$. On the other hand, Lemma 2.15 gives that $\omega$-almost surely $(x_{1,k},z_{1,k})$ is $i$-uncompressed, so $j\neq i$ - a contradiction.
We conclude the existence of a function $d_{0}$ as declared in the beginning of step 1: for all $d\geq d_{0}(x)$, all $d$-separated $i$-horizontal pairs in the ball $B:=B\big(x,99d\big)$ are simultaneously $j$-compressed or simultaneously $j$-uncompressed.
Step 22
We now show that we can choose $d_{0}\in O(u)$. We start by taking, for each $x\in X$, the infimal $\tilde{d}_{0}(x)$ such that for all $d\geq\tilde{d}_{0}(x)$, all $i$-horizontal $d$-separated pairs in $B(x,99d)$ are $j$-compatible. Next define the radial nondecreasing function $d_{0}(|x|):=\sup_{y:|y|\leq|x|}\tilde{d}_{0}(y)$.
A-priori $d_{0}:\mathbf{R}_{\geq 0}\rightarrow\mathbf{R}_{\geq 0}\cup\{\infty\}$, and we need to show it is in fact real valued (this is trivial if the spaces are uniformly discrete). Assume towards contradiction that $d_{0}$ is not bounded in some ball $B:=B(x_{0},R)$, and get points $x_{k}\in B$ and real numbers $r_{k}\rightarrow\infty$ such that the balls $B(x_{k},99r_{k})$ contain two $i$-horizontal $r_{k}$-separated pairs $(x_{1,k},z_{1,k}),(x_{2,k},z_{2,k})$ that are $j$-incompatible - say that $(x_{1,k},z_{1,k})$ are $j$-compressed while $(x_{2,k},z_{2,k})$ are $j$-uncompressed. We consider the cone $X_{\omega}:=(X,x_{0},r_{k})$, that is $u$-admissible. Since $x_{k}\in B$ we have $d(x_{k},x_{0})\leq R$, so $x_{\omega,\alpha}:=[(x_{\alpha,k})],z_{\omega,\alpha}:=[(z_{\alpha,k})]\in X_{\omega}$. By Lemma 2.15 we see that $j$ cannot equal $i$ because of the assumption on $(x_{1,k},z_{1,k})$, and $j$ must equal $i$ because of the assumption on $(x_{2,k},z_{2,k})$ - a contradiction.
The function $d_{0}$ is therefore a radial, real valued, bounded near 0, and nondecreasing. We show it is $O(u)$. Assume towards contradiction that there is a sequence $x_{k}\rightarrow\infty$ so that $\lim_{\omega}\frac{u(x_{k})}{d_{0}(x_{k})}=0$. By the construction of $d_{0}$ we may take $x_{k}^{\prime}$ such that $|x_{k}^{\prime}|\leq|x_{k}|$, $d_{0}(x_{k}^{\prime})=d_{0}(x_{k})$, and for which $B\big(x_{k}^{\prime},99d_{0}(x_{k}^{\prime})\big)$ contains two $\big(d_{0}(x_{k}^{\prime})-1\big)$-separated $i$-horizontal pairs $(x_{1,k},z_{1,k}),(x_{2,k},z_{2,k})$ that are $j$-incompatible. Since $u$ is nondecreasing, the assumption $\lim_{\omega}\frac{u(x_{k})}{d_{0}(x_{k})}=0$ implies that the cone with moving base points $X_{\omega}:=\operatorname{Cone}_{\omega}\big(X,x_{k}^{\prime},d_{0}(x_{k}^{\prime})\big)$ is a $u$-admissible asymptotic cone. Since $x_{\alpha,k},z_{\alpha,k}\in B\big(x_{k}^{\prime},99d_{0}(x_{k})\big)$, we see that $x_{\omega,\alpha}:=[(x_{\alpha,k})],z_{\omega,\alpha}:=[(z_{\alpha,k})]\in X_{\omega}$. Once again, Lemma 2.15 yields a contradiction.
Step 33
We have established the following: there is $d_{0}\in O(u)$ such that for every $x\in X$ and every $d\geq d_{0}(x)$, the property of whether a $d$-separated $i$-horizontal pair inside $B=B(x,99d)$ is $j$-uncompressed or $j$-compressed depends only on $x$, $i$, $j$ and $d$, and not on the specific pair (in particular such a pair cannot be $j$-semi-compressed). We call it the $j$-type of $B$. Since $d_{0}\in O(u)$ and nondecreasing we may assume that $d_{0}$ is also subadditive. We can right away conclude that there are no $d_{0}$-separated $i$-horizontal $j$-semi-compressed pairs at all in $X$: such a pair $(x,x^{\prime})$ must admit $d(x,x^{\prime})\geq d_{0}(x)$ and therefore for $d=d(x,x^{\prime})\geq d_{0}(x)$, $x,x^{\prime}$ are $d$-separated and $x^{\prime}\in B(x,99d)$. Step $1$ shows that $(x,x^{\prime})$ is not $j$-semi-compressed.
Our next goal is to remove any restriction on the radius of the balls $B(x,99d)$. Fix $x\in X$. First observe that the $j$-type of $d_{0}$-separated $i$-horizontal pairs which involve $x$ depend only on $x$, as they all have the same type as the type of the ball $B:=B\big(x,99d_{0}(x)\big)$. Explicitly, let $(x,x^{\prime})$ be a $d_{0}$-separated $i$-horizontal pair. If $x^{\prime}\in B$, $(x,x^{\prime})$ has the same $j$-type as $B$. If $x\notin B$, we may connect $x$ and $x^{\prime}$ by a geodesic $\gamma:=[x,x^{\prime}]$, and subdivide this geodesic into subintervals for which the endpoint have the same type as that of $B$. Precisely, the pair $\Big(x,\gamma\big(98d_{0}(x)\big)\Big)$ is $d>d_{0}(x)$-separated and $i$-horizontal, so it has the $j$-type of $B$. Therefore for $R_{1}:=98d_{0}(x)$, the ball $B_{1}:=B(x,99R_{1})$ has the $j$-type of $B$. We can continue to enlarge the radius of the balls until we find one which contains $x^{\prime}$, so indeed $(x,x^{\prime})$ has the $j$-type of $B$, which is what we were set out to prove
[…]
For clarity of presentation we make the following notations.
Definition 2.18.
For $a,b\in\mathbf{R}$, denote $a\lesssim b$ if there is $M>0$ such that $a\leq Mb$, and $a\sim b$ if both $a\lesssim b$ and $b\lesssim a$ hold. Let $(a_{k})_{k},(b_{k})_{k}$ be sequences of real numbers. We denote
@@
    $a_{k}\lesssim b_{k}$ if there is $M>0$ independent of $k$ such that $a_{k}\leq Mb_{k}$ for all large enough $k$, and $a_{k}\sim b_{k}$ if both $a_{k}\lesssim b_{k}$ and $b_{k}\lesssim a_{k}$ hold.
@@
    $a_{k}\lesssim_{\omega}b_{k}$ if there is $M>0$ independent of $k$ such that $a_{k}\leq Mb_{k}$ $\omega$-almost-surely, and $a_{k}\sim_{\omega}b_{k}$ if both $a_{k}\lesssim_{\omega}b_{k}$ and $b_{k}\lesssim_{\omega}a_{k}$ hold.
We will repeatedly and implicitly use the following easy facts for an $O(v)$-bilipschitz embedding:
  1. (1)
    If $(z_{k})$ is a sequence in $Z$ with $|z_{k}|\rightarrow\infty$, then $|z_{k}|\sim|f(z_{k})|$.
  2. (2)
    If $(w_{k})$ is a sequence in $W$ with $|w_{k}|\rightarrow\infty$, and $|z_{k}|$ are such that $d_{W}\big(f(z_{k}),w_{k}\big)\leq v(w_{k})$, then $|z_{k}|\sim|w_{k}|$. Such $z_{k}$ exist if $f$ is an $O(v)$-equivalence.
Proof of Theorem C.
@@
From Lemma 2.16 we conclude that for every $i$ there is a unique $j$ for which the projection on $Y_{j}$ is uncompressed for all $d_{0}$-separated $i$-horizontal pairs. To ease notation we assume $j=i$ is that unique index, and fix from now on $i=j=1$. The proof is identical for all other $i$.
@@
$L_{1}^{-1}d(z,z^{\prime})\leqslant d\Big(\pi_{Y_{1}}\big(\phi(z,\widehat{x})\big),\pi_{Y_{1}}\big(\phi(z^{\prime},\widehat{x})\big)\Big)\leqslant L_{1}d(z,z^{\prime}).$